Circular No. 8977

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2008 S1 (McNAUGHT) R. H. McNaught reports his discovery of a comet (discovery position

tabulated below) with a 15′′ circular coma and possible extension in p.a. ∼ 30^{o} on CCD images obtained with the 0.5-m Uppsala Schmidt telescope; additional stacked images from Sept. 18.38 UT show a 30′′ coma extended to the north and a condensation with FHWM = 5′′. Following posting on the ‘NEOCP’ webpage, other CCD observers have also reported cometary appearance. D. Mayes and J. Young (Table Mountain 0.6-m reflector, Sept. 18.1; low altitude) report a round coma with an 8′′-diameter central condensation and a possible thin tail ≈ 20′′ long in p.a. 55^{o}–60^{o}. E. Guido, G. Sostero, and P. Camilleri (0.25-m reflector at Moorook, Australia, remotely, Sept. 18.5; 31 co-added exposures) find a coma diameter of ∼ 12′′. A. Herring measures a 9′′.9 tail in p.a. 59^{o}.4 on exposures taken by R. Holmes (Charleston, IL, U.S.A., 0.61-m astrograph, Sept. 18.05–18.06).

2008 UT  α_{2000}  δ_{2000}  Mag.
Sept.17.40807  16^{h}21^{m}33.^{s}37  −24^{o}29′41′′.3  16.6

The available astrometry, preliminary parabolic orbital elements (T = 2008 Sept. 21.089 TT, q = 1.35320 AU, ω = 192^{o}.082, Ω = 104^{o}.558, i = 17^{o}.974, equinox 2000.0), and an ephemeris appear on MPEC 2008-S10.

S/2008 (35107) 1 W. J. Merline, Southwest Research Institute (SwRI); A. R. Conrad,

W. M. Keck Observatory (WMKO); J. D. Drummond, Starfire Optical Range, AFRL; P. M. Tamblyn, Binary Astronomy, Dillon, CO, and SwRI; C. Dumas and B. Carry, European Southern Observatory; R. D. Campbell and R. W. Goodrich, WMKO; C. R. Chapman, SwRI; and W. M. Owen, Jet Propulsion Laboratory, report the first-ever near-infrared imaging of a close-approaching binary minor planet: on Aug. 9 UT, over a span of ∼ 1 hr, they obtained J-, H-, and K_{p}-band images — using the 10-m Keck II Telescope (+ NIRC2/AO adaptive-optics system) on Mauna Kea — of clearly-separated components of (35107) 1991 VH, which was first suspected as a binary by Pravec et al. (it IAUC 6607). On Aug. 9.236 the satellite, designated S/2008 (35107) 1, was 0′′.08 (projected separation 3.1 km) in p.a. 105^{o} from the primary (the smallest angular and physical separations yet for a binary minor planet using adaptive optics). The brightness difference was ΔK_{p} ≈ 2.0 mag. The observed separation and size ratio are consistent with the parameters derived by Pravec et al. (2006, Icarus 181, 63).

2008 September 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green